OpenStudy (anonymous):

verify the idenity? (acost-bsint)^2+(asint+bcost)^2=a^2+b^2

OpenStudy (zehanz):

First, you have to know how to expand forms like (a+b)² and (a-b)². Can you do that?

OpenStudy (anonymous):

(acost + bsint) (acost-bsint) for the (a-b)^2?

OpenStudy (zehanz):

No, I mean (acost-bsint)² expands just like (p-q)², where p=acost and q=bsint, So exapnd it first, then handle the other one.

OpenStudy (zehanz):

Let me help you: (p-q)²=p²-2pq+q², so (acost-bsint)²=...

OpenStudy (anonymous):

(acost-bsint)^2= (acost^2-2(acost)(bsint)+bsint^2?

OpenStudy (zehanz):

It is \(a^2\cos^2t-2ab\sin t \cos t + b^2 \sin^2 t\) Then, (asint+bcost)² expands to: \(a^2\sin^2 t+2ab\sin t \cos t +b^2 \cos^2 t \)

OpenStudy (zehanz):

What happens if you add these two beasts?

OpenStudy (anonymous):

turns into the right side? the middle section cancels, right?

OpenStudy (zehanz):

Yes, they cancel, so we are left with: \(a^2(\cos^2t+\sin^2t)+b^2(\sin^2t+\cos^2t)\) What happens with the stuff between the brackets?

OpenStudy (anonymous):

turns into 1, due to the Pythagorean identity. You made it look so easy, hopefully it sticks in my head.

OpenStudy (zehanz):

OK, next time you will know what to do!

OpenStudy (anonymous):

Thank you, much appreciated!